OpenStudy (anonymous):

You roll a standard number cube. Find P(3 or 4).

OpenStudy (anonymous):

By a standard number cube, do you mean a dice? If so then the probabilty of rolling a number, x is as follows: P(1)=1/6 as there is one face on a die that has a number 1, and six faces it can land on P(2)=1/6 P(3)=1/6 etc etc So the probability of rolling P(3) = 1/6 and P(4)=1/6 then the P(3) or P(4) = P(3 or 4) = ?

OpenStudy (anonymous):

1/3 = P(3or4)